OpenStudy (anonymous):

anti-derivative of SQRT(1+(9x- (x^3/3))^2)

OpenStudy (amistre64):

how you wanna do this lol the easy way or the hard way

OpenStudy (anonymous):

lets go with easy because this is the part of problems i can never do

OpenStudy (amistre64):

\[\sqrt{1+(9x-\frac{x^3}{3})^2}\] lets start by making sure we got the right problem is this it?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

well; its gonna be a trig sub that turns to int(sec(x)) dx inthe end lol

OpenStudy (anonymous):

omg i hate calculus lol

OpenStudy (amistre64):

lol...trig sub just cleans it up thats all

OpenStudy (anonymous):

yea but i integrated this on my calculator and i still got the wrong answer

OpenStudy (amistre64):

\[\int_{} \sqrt{1+tan^2(u)}du\] ;\[u= \frac{27x -x^3}{3}\]

OpenStudy (amistre64):

du = 27-3x^2/3 = 9-x^2 dx; maybe that aint a good idea lol

OpenStudy (amistre64):

but it shows me another idea lol

OpenStudy (anonymous):

yea im not so good with substitution either

OpenStudy (amistre64):

u = 9x-x^3 ; du = 9-3x^2 dx {S} sqrt(1+(u/3)^2) du -------------- 9-x^2 maybe?

OpenStudy (amistre64):

let me stare blindly at it for a few moments and let my thoughts congeal lol

OpenStudy (anonymous):

ok im just trying to follow the process here

OpenStudy (amistre64):

we might be better of expanding and combining everything under the radical into one shot

OpenStudy (anonymous):

sure i couldnt go anymore wrong with this problem lol

OpenStudy (amistre64):

\[\frac{\sqrt{x^6 + 729x^2 -54x^4 +9}}{3}\] is what i get with that ...maybe

OpenStudy (anonymous):

yea i got something completely off but u r def prolly right lol

OpenStudy (amistre64):

now we can factor under the radical to simply fy things

OpenStudy (amistre64):

sooo...how should we factor it :) all polynomials higher then degrees 2 can factor to linears and or irreducibla quads

OpenStudy (anonymous):

i honestly have no idea lol

OpenStudy (amistre64):

ima try some synthetic stuff and hope i get lucky

OpenStudy (anonymous):

lol ok im attempting to do other problems..im going to fail this quiz tomorrow lol

OpenStudy (amistre64):

well, the good news is; wolfram couldnt even figure it out lol

OpenStudy (anonymous):

whos that? some really smart kid im guessing :)

OpenStudy (amistre64):

www.wolframalpa.com type in: int(sqrt(1+(9x-(x^3)/3)^2))dx

OpenStudy (anonymous):

o gotcha lol..ok well maybe i went wrong somewhere..the question is find the length of the arch of the parabola y=9-x^2 that lies above the x-axis

OpenStudy (amistre64):

wolframalpha.com

OpenStudy (amistre64):

.... why you wanna keep secrets from me; i cant read your mind

OpenStudy (amistre64):

you had the wrong stuff inputed lol

OpenStudy (anonymous):

i did?? well the equation we use is SQRT(1+(dy/dx)^2

OpenStudy (amistre64):

\[\int\limits_{} \sqrt{1 + [f'x]^2} dx\]

OpenStudy (amistre64):

and is the derivative of 9-x^2 = to the integral of 9-x^2?

OpenStudy (anonymous):

yea and the derivative of 9-x^2 is 9x-(x^3/3) right?

OpenStudy (amistre64):

dy/dx = -2x

OpenStudy (anonymous):

ooo i did the anti derivative didnt i ..i do that all the time because we keep having to switch from derivativeto anti derivative lol

OpenStudy (amistre64):

anti derivative is a useless term; later chapters call them what they are inegrals lol

OpenStudy (amistre64):

integrals

OpenStudy (anonymous):

yea that what we do when we use integrals

OpenStudy (amistre64):

{S} sqrt(1+4x^2) dx ; [a,b] whatever your interval s its up there some where

OpenStudy (anonymous):

ok now i got it..wow i am retarded lol i do that a lot to

OpenStudy (anonymous):

so now how to integrate SQRT(1+4x^2)

OpenStudy (amistre64):

easier than you would that other monstrocity lol

OpenStudy (anonymous):

i got 2/3(1+4x^2)^(3/2) and then do you also multiply that by the derivative of the inside??

OpenStudy (amistre64):

sqrt(1+tan^2(u)) sec^2(u) du ------------------------- 4 sec(u). sec^2(u) du ---------------- 4 sec^3(u) du ---------- 4

OpenStudy (amistre64):

got that outta whack a little

OpenStudy (anonymous):

ok i g2g i got soccer..but thanks for the help :)